### v2

# Least multivariate Chebyshev polynomials on diagonally-determined domains

Abstract.
We consider a new multivariate generalization of the classical monic (univariate) Chebyshev polynomial that minimizes the uniform norm on the interval $[-1,1]$. Let $\Pi^{*}_{n}$ be the subset of polynomials of degree at most $n$ in $d$ variables, whose homogeneous part of degree $n$ has coefficients summing up to $1$. The problem is determining a polynomial in $\Pi^{*}_{n}$ with the smallest uniform norm on a domain $\Omega$, which we call a least Chebyshev polynomial (associated with $\Omega$). Our main result solves the problem for $\Omega$ belonging to a non-trivial class of domains that we call diagonally-determined, and establishes the remarkable result that a least Chebyshev polynomial can be given via the classical, univariate, Chebyshev polynomial. In particular, the solution can be independent of the dimension. Diagonally-determined domains include centered balls in $\mathbb{R}^{d}$ in any norm, but can be non-convex and highly irregular. We also introduce a computational procedure, based on semidefinite programming hierarchies, to detect if a given semi-algebraic set is diagonally-determined.
Key words and phrases:
Least polynomial, Chebyshev polynomial, best approximation, extremal signature
2020 Mathematics Subject Classification:
Primary: 41A10, 41A63.
1. Introduction
Among its numerous properties, the Chebyshev polynomial $T_{n}(x)=\cos(n\arccos x)$ provides a solution for the best approximation to the monomial $x^{n}$ on the interval $[-1,1]$ in the uniform norm. More precisely, the polynomial
$q_{n}^{*}(x)=x^{n}-2^{1-n}T_{n}(x)$
of degree $n-1$ ($n\in\mathbb{N}$) is the best polynomial of approximation to $x^{n}$ on $[-1,1]$; that is
(1.1) $q_{n}^{*}\,=\,\arg\min_{q\in\Pi_{n-1}}\,\sup_{x\in[-1,1]}|x^{n}-q(x)|\,,$
with $\Pi_{n-1}$ being the vector space of univariate polynomials of degree at most $n-1$. In other words, the monic Chebyshev polynomial $x^{n}-q^{*}_{n}$ is the least polynomial in the sense that it has the least uniform norm among all monic polynomials of degree $n$.
There have been multiple extensions of Chebyshev polynomials to multivariate settings from different angles. From the point of view of approximation, an immediate generalization is finding the best approximation to monomials [1, 2, 5, 3, 9, 7, 11, 15, 16, 17]. Namely, for $d>1$, ${\alpha}\in{\mathbb{N}}^{d}$ and $|{\alpha}|\coloneqq\sum_{i=1}^{d}\alpha_{i}=n$, we consider the problem
(1.2) $\min_{q\in\Pi^{d}_{n-1}}\,\sup_{{\mathbf{x}}\in\Omega}\,|{\mathbf{x}}^{{\alpha}}-q({\mathbf{x}})|\,,$
where $\Pi^{d}_{n-1}$ denotes the real polynomials of total degree at most $n-1$ in the variables ${\mathbf{x}}=(x_{1},\ldots,x_{d})$, and where we define the monomial ${\mathbf{x}}^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\ldots x_{d}^{\alpha_{d}}$. Since we will deal with monomials of degree $n$ throughout the paper, we introduce the notation
${\mathbb{N}}^{d}_{n}\coloneqq\left\{\alpha\in({\mathbb{N}}_{0})^{d}\;:\;|\alpha|=n\right\},$
where ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$ denotes the nonnegative integers. Thus ${\mathbb{N}}^{d}_{n}$ indexes all monomials of degree exactly $n$ in $d$ variables.
Problem (1.2) can be regarded as a natural multivariate generalization of (1.1), where $\Omega\subset{\mathbb{R}}^{d}$ is a subset of ${\mathbb{R}}^{d}$. While the interval $[-1,1]$ is a prototype of a compact connected set of the real line, there is no ‘prototype’ in higher dimensions for such a set. In the literature, this problem has been studied primarily on a few special regular domains. In two variables, the problem (1.2) is solved for the square, the disk, and the isosceles right triangle [5, 3, 7, 11, 15]. While the solution on the square can be extended to the cube for $d>2$, the problem is solved only for a few cases, mostly monomials of lower degrees, on the ball and the simplex [1, 2, 16]. Moreover, the existing examples indicate an increasing complexity, so much so that it does not appear possible to find an analytic solution even for these regular domains.
Recently, in [4] we have proposed to investigate (1.2) for various choices of $\Omega\subset{\mathbb{R}}^{d}$ and ${\alpha}\in{\mathbb{N}}^{d}_{n}$ by combining analytical tools with numerical tools from optimization (and notably the so-called moment-SOS hierarchy). During this study, we have encountered an optimization problem that has initiated a change of view: Namely, instead of studying the best polynomial of approximation to monomials, we can study the least polynomial instead. While the two concepts are identical in one variable, they can be quite different in higher dimensions, as seen from the definition below.
Definition 1.1.
Let $\Pi_{n}^{d}$ denote the space of polynomials of total degree at most $n$ in $d$ variables, and $\Pi^{*}_{n}$ the subset of $\Pi_{n}^{d}$ that consists of polynomials of the form
${\mathbf{x}}\mapsto P({\mathbf{x}})\,\coloneqq\,\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}\,{\mathbf{x}}^{\alpha}+Q({\mathbf{x}})\,\text{ with }\quad\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1\,\text{and }\>Q\in\Pi^{d}_{n-1}.$
Let $\Omega$ be a domain in ${\mathbb{R}}^{d}$. We consider the optimization problem
(1.3) $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega},\quad\hbox{where}\quad\|P\|_{\Omega}\coloneqq\sup_{{\mathbf{x}}\in\Omega}\,|P({\mathbf{x}})|\,.$
If it exists, we call a minimizer $P^{*}\in\Pi^{*}_{n}$ of (1.3) a least Chebyshev polynomial of degree $n$ on the domain $\Omega$.
For $d=1$, there is only one monomial of degree $n$. In the case $d>1$, every element of $\Pi^{*}_{n}$ is ‘monic’ and the monomial ${\mathbf{x}}^{\alpha}$ in (1.1) is only one among many possible choices in $\Pi^{*}_{n}$. However, rather than approximating a fixed monomial by polynomials of lower degree, the problem (1.3) requires finding a polynomial that has the least norm among all polynomials in $\Pi_{n}^{*}$. As far as we are aware, this polynomial has not been considered in the literature.
Contribution
The main purpose of this paper is to report our findings on the optimization problem (1.3). It turns out, much to our surprise, that the problem (1.3) can be solved analytically for a fairly general family of domains $\Omega$ in ${\mathbb{R}}^{d}$ for all $d\geq 2$. This family of domains will be referred to as diagonally-determined.
Organization of this paper
The paper is organized as follows. We start by defining and describing several examples of diagonally-determined domains in Section 2. Our main results are presented in Section 3, where we describe the least polynomial for a diagonally-determined domain. We discuss the dual problem of (1.3) in Section 4 and show that it too has a closed-form solution in the case of diagonally-determined sets. We also rephrase this result for the dual problem in the framework of extremal signatures. Finally, we consider the problem of deciding whether a given semi-algebraic set is diagonally determined in Section 5, and propose an algorithmic approach based on semidefinite programming hierarchies. We conclude the paper with a discussion of related problems and open questions in Section 6.
2. Diagonally-determined domains
We define the diagonal of a domain $\Omega\subset\mathbb{R}^{d}$ as the set ${\rm diag}(\Omega)\coloneqq\{t\in{\mathbb{R}}:t\mathbf{1}\in\Omega\}$, where $\mathbf{1}$ denotes the all-ones vector in $\mathbb{R}^{d}$. For two vectors ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{d}$, we use $\langle{\mathbf{x}},{\mathbf{y}}\rangle$ for their standard inner product.
Definition 2.1.
We call a set $\Omega\subset\mathbb{R}^{d}$ diagonally-determined if the following two conditions hold:
  1. (1)
    The diagonal of $\Omega$ is an interval, say ${\rm diag}(\Omega)=[a,b]$, and
  2. (2)
    there exists ${\mathbf{v}}\in{\mathbb{R}}^{d}$ such that $\langle{\mathbf{v}},\mathbf{1}\rangle=1$ and $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in[a,b]$ for all ${\mathbf{x}}\in\Omega$.
Importantly, Definition 2.1 covers domains that could be non-convex, non-compact, and even highly irregular. Figure 1 gives an example of a non-convex diagonally-determined set in ${\mathbb{R}}^{2}$.
$x_{1}$$x_{2}$001111222233334444111122223333
Figure 1. Example of a diagonally-determined, non-convex set. Here, $\mbox{diag}(\Omega)=[1,3]$, and ${\mathbf{v}}=(1,0)$.
As a first observation, a dilation of a diagonally-determined set is again diagonally-determined.
Lemma 2.2.
If $\Omega\subset\mathbb{R}^{d}$ is diagonally-determined with ${\rm diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$, then, for any $r>0$, the dilation
$r\Omega\coloneqq\{r{\mathbf{x}}\;:\;{\mathbf{x}}\in\Omega\}$
is also diagonally-determined with ${\rm diag}(r\Omega)=[ra,rb]$ and vector ${\mathbf{v}}$.
Proof.
The proof is an immediate consequence of Definition 2.1. ∎
We will now show that balls in ${\mathbb{R}}^{d}$ (in any norm), centered at the origin, are examples of diagonally-determined sets.
Proposition 2.3.
Consider a ball of radius $r$ in ${\mathbb{R}}^{d}$ centered at the origin:
$\Omega\coloneqq\{{\mathbf{x}}\in{\mathbb{R}}^{d}\;:\;\|{\mathbf{x}}\|\leq r\},$
[…]
$\|{\mathbf{x}}\|\leq\||{\mathbf{x}}|\|\quad\forall{\mathbf{x}}\in\mathbb{R}^{d},$
where $|{\mathbf{x}}|=(|x_{1}|,\ldots,|x_{d}|)$, then the vector ${\mathbf{v}}$ may be assumed to be entrywise-nonnegative, denoted by  ${\mathbf{v}}\geq\mathbf{0}$.
Proof.
We prove the statement for unit balls; the required result then follows from Lemma 2.2.
[…]
$\|{\mathbf{u}}\|_{*}=\sup\{\langle{\mathbf{u}},{\mathbf{x}}\rangle\;:\;\|{\mathbf{x}}\|=1\},\quad{\mathbf{u}}\in{\mathbb{R}}^{d}.$
Let ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, where ${\mathbf{u}}\in{\mathbb{R}}^{d}$ is ‘dual’ to $\mathbf{1}$, in the sense that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$. We then have ${\rm diag}(\Omega)=[a,b]$ with $a=-1/\|\mathbf{1}\|$ and $b=1/\|\mathbf{1}\|$, $\langle{\mathbf{u}}/\|\mathbf{1}\|,\mathbf{1}\rangle=1$, and $|\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle|\leq(1/\|\mathbf{1}\|)\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq 1/\|\mathbf{1}\|=b$, so that $\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle\in[a,b]$ for all ${\mathbf{x}}\in\Omega$.
[…]
$\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\quad\forall{\mathbf{u}}\in\mathbb{R}^{d}.$
Thus, the vector ${\mathbf{u}}$ ‘dual’ to $\mathbf{1}$ in the above construction is component-wise nonnegative. Indeed, if not, then
$1=\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\geq\langle|{\mathbf{u}}|,\mathbf{1}/\|\mathbf{1}\|\rangle>\langle{\mathbf{u}},\mathbf{1}/\|\mathbf{1}\|\rangle=1,$
a contradiction. Subsequently, since ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, we have ${\mathbf{v}}\geq\mathbf{0}$, as required. ∎
By Definition 2.1, we have the following immediate corollary.
Corollary 2.4.
Assume $\Omega\subset{\mathbb{R}}^{d}$ is a subset of a ball in ${\mathbb{R}}^{d}$ in any norm, centered at the origin, and that the diagonal of $\Omega$ coincides with the diagonal of the ball. Then $\Omega$ is a diagonally-determined set.
To illustrate this corollary, the example in Figure 2 shows a non-convex subset of a unit (Euclidean) ball, that has the same diagonal as the ball.
[…]
Figure 2. Example of a diagonally-determined, non-convex set. Here, $\mbox{diag}(\Omega)=\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$, and ${\mathbf{v}}=\left(\frac{1}{2},\frac{1}{2}\right)$.
Example 2.5.
It is insightful to consider the specific example of the unit ball of the owl norm (ordered weighted $\ell_{1}$-norm) defined relative to weights $w_{1}\geq w_{2}\geq\cdots\geq w_{d}\geq 0$ by
$\|{\mathbf{x}}\|_{\sf owl}=\sum_{i=1}^{d}w_{i}x^{*}_{i},$
with $(x_{1}^{*},\ldots,x_{d}^{*})$ being the nonincreasing rearrangement of $(|x_{1}|,\ldots,|x_{d}|)$. Note that $\|\mathbf{1}\|_{\sf owl}=\sum_{i=1}^{d}w_{i}\eqqcolon W$, so that $b=1/W$. Letting ${\mathbf{w}}=(w_{1},\ldots,w_{d})$, we set ${\mathbf{v}}={\mathbf{w}}/W$, since $\langle{\mathbf{w}}/W,\mathbf{1}\rangle=(\sum_{i=1}^{d}w_{i})/W=1$ and, for $\|{\mathbf{x}}\|_{\rm owl}\leq 1$,
$|\langle{\mathbf{w}}/W,{\mathbf{x}}\rangle|$ $=$ $\left|\sum_{i=1}^{n}w_{i}x_{i}\right|/W$
$\leq$ $\left(\sum_{i=1}^{n}w_{i}|x_{i}|\right)/W$
$\leq$ $\left(\sum_{i=1}^{n}w_{i}x^{*}_{i}\right)/W$
$\leq$ $1/W=b.$
The intersection of certain balls in ${\mathbb{R}}^{d}$ with the nonnegative orthant in ${\mathbb{R}}^{d}$ is also diagonally-determined, as the next result shows.
Proposition 2.6.
[…]
Further assume $\Omega\subset{\mathbb{R}}^{d}$ has diagonal $\mathrm{diag}(\Omega)=[0,r/\|\mathbf{1}\|]$ for some $r>0$, and, for all ${\mathbf{x}}\in\Omega$, ${\mathbf{x}}\geq\mathbf{0}$ and $\|{\mathbf{x}}\|\leq r$. Then $\Omega$ is a diagonally-determined set with vector ${\mathbf{v}}=\mathbf{u}/\|\mathbf{1}\|$.
Proof.
We again prove the statement for the case $r=1$; the result for general $r>0$ then follows from Lemma 2.2. Setting ${\mathbf{v}}=\mathbf{u}/\|\mathbf{1}\|$, one has $\langle{\mathbf{v}},\mathbf{1}\rangle=1$. Moreover, for ${\mathbf{x}}\in\Omega$, $\langle{\mathbf{v}},{\mathbf{x}}\rangle\geq 0$, since ${\mathbf{x}},{\mathbf{v}}\geq 0$, and
$\langle{\mathbf{v}},{\mathbf{x}}\rangle=\frac{1}{\|\mathbf{1}\|}\langle\mathbf{u},{\mathbf{x}}\rangle\leq\frac{1}{\|\mathbf{1}\|}\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq\frac{1}{\|\mathbf{1}\|},$
so that $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in\mathrm{diag}(\Omega)$ for all ${\mathbf{x}}\in\Omega$. ∎
We note again that the assumption on the norm in Proposition 2.6 is met by, for example, all monotone norms, as shown in the proof of Proposition 2.3.
As a simple corollary of the proposition, the simplex
$\Omega=\left\{{\mathbf{x}}\in{\mathbb{R}}^{d}\;:\;{\mathbf{x}}\geq\mathbf{0},\;\sum_{i=1}^{d}x_{i}\leq 1\right\},$
is diagonally-determined, since it is the intersection of the unit $\ell^{1}$-ball with the nonnegative orthant. See Figure 3 for an illustrative example of Proposition 2.6.
[…]
Figure 3. Example to illustrate Proposition 2.6. Here, $\mbox{diag}(\Omega)=[0,b]$, with $b\approx 1.35$. Moreover, $\Omega$ is contained in the $\ell^{1}$-ball of radius $r=b{d}\approx 2.7$, intersected with the nonnegative quadrant.
Finally, the diagonally-determined property is not invariant under translation in general. The exception is a translation by a constant multiple of the all-ones vector. We state this observation as a lemma for later use. The proof again follows immediately from Definition 2.1 and is omitted.
Lemma 2.7.
Fix $\alpha\in\mathbb{R}$. Then a given $\Omega\subset\mathbb{R}^{d}$ is diagonally determined with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$ if and only if its translation $\Omega+\alpha\mathbf{1}$ is diagonally determined with diagonal $[a+\alpha,b+\alpha]$ and vector ${\mathbf{v}}$.
3. Least Chebyshev polynomial for a diagonally-determined domain
In this section, and with $\Omega$ being a domain in ${\mathbb{R}}^{d}$, we consider the problem of finding a least polynomial $P^{*}\in\Pi_{n}^{*}$ such that
$\|P^{*}\|_{\Omega}=\inf\left\{\|P\|_{\Omega}:P\in\Pi_{n}^{*}\right\},$
for the special case where $\Omega$ is diagonally-determined.
Theorem 3.1.
Let $\Omega$ be diagonally-determined with vector ${\mathbf{v}}$ and $\mathrm{diag}(\Omega)=[a,b]$. Then
(3.1) $\inf\left\{\|P\|_{\Omega}:P\in\Pi_{n}^{*}\right\}=\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}\,,$
and the infimum is attained by the polynomial
(3.2) ${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})\,\coloneqq\,\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(-1+2\,\frac{\langle{\mathbf{v}},{\mathbf{x}}\rangle-a}{b-a}\right).$
Proof.
Let $P\in\Pi_{n}^{*}$, say
$P({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}\,{\mathbf{x}}^{\alpha}+Q({\mathbf{x}})\,\quad\text{with }\quad\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1\,\text{and }\>Q\in\Pi^{d}_{n-1}.$
Then $P(t,\ldots,t)=t^{n}+q_{n-1}(t)$, where $q_{n-1}(t)=Q(t,\ldots,t)$ is a univariate polynomial of degree at most $n-1$ in the single variable $t$. Hence,
(3.3) $\|P\|_{\Omega}\geq\max_{a\leq t\leq b}\left|t^{n}+q_{n-1}(t)\right|\geq\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}\,,$
where we have used a classical result in one variable, and equality in the last inequality is attained by choosing $q_{n-1}$ such that $t^{n}+q_{n-1}(t)$ is the rescaled Chebyshev polynomial in the right-hand side of (3.2).
Moreover, as $\Omega$ is diagonally-determined, then for all ${\mathbf{x}}\in\Omega$, one has $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in[a,b]$, and therefore
$\|P^{*}\|_{\Omega}=\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}\max_{{\mathbf{x}}\in\Omega}\left|T_{n}\left(-1+2\frac{\langle{\mathbf{v}},{\mathbf{x}}\rangle-a}{b-a}\right)\right|\leq\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}\,.$
Finally, we need to show that $P^{*}\in\Pi_{n}^{*}$. First note that $t\mapsto P^{*}(t\mathbf{1})$ is a monic univariate polynomial in $t$, since $\langle{\mathbf{v}},\mathbf{1}\rangle=1$. On the other hand, we may write $P^{*}$ in the form
$P^{*}({\mathbf{x}})=\sum_{\alpha\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}\,{\mathbf{x}}^{\alpha}+Q^{*}({\mathbf{x}})\,\text{ with }Q^{*}\in\Pi^{d}_{n-1},$
so that
$P^{*}(t\mathbf{1})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}t^{n}+Q^{*}(t\mathbf{1}).$
Thus $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}=1$, and therefore $P^{*}\in\Pi_{n}^{*}$, as required. ∎
The minimal value in (3.1) does not depend explicitly on the dimension $d$, but the diagonal $[a,b]$ may. For example, if $\Omega$ is the unit Euclidean ball in ${\mathbb{R}}^{d}$, then $[a,b]=[-1/\sqrt{d},1/\sqrt{d}]$.
Example 3.2.
For the cube $\Omega=[-1,1]^{d}$, Theorem 3.1 shows that $\|P^{*}\|_{\Omega}=2^{-n+1}$. In contrast, for every ${\alpha}\in{\mathbb{N}}^{d}$ with $|\alpha|=n$, we have
$\inf_{P\in\Pi_{n-1}^{d}}\|{\mathbf{x}}^{\alpha}-P\|_{[-1,1]^{d}}=2^{-n+d},$
as shown in [18] (see also [15]), which depends on the dimension $d$.
Since an optimal solution of problem (1.3) is given in terms of the univariate Chebyshev polynomial $T_{n}$ in Theorem 3.1, the reader may wonder if this optimal solution is in fact unique. This turns out to be not the case, as the next example shows.
Example 3.3.
[…]
$\|P^{*}\|_{\triangle^{d}}=\inf\left\{\|P\|_{\triangle^{d}}:P\in\Pi_{n}^{*}\right\}=\left(\frac{1}{2}\right)^{n}\frac{1}{2^{n-1}}=2^{-2n+1}$
is independent of the dimension $d$. Moreover, one has $P^{*}({\mathbf{x}})=2^{-2n+1}T_{n}(2x_{1}-1)$, which is univariate. Another valid choice for ${\mathbf{v}}$ is ${\mathbf{v}}=(0,1,0,\ldots,0)$ that leads to a different least Chebyshev polynomial, namely ${\mathbf{x}}\mapsto 2^{-2n+1}T_{n}(2x_{2}-1)$.
Example 3.4.
Returning to Example 2.5, where $\Omega$ is the unit ball for the owl norm, we have $[a,b]=[-1/W,1/W]$ and ${\mathbf{v}}={\mathbf{w}}/W$, so that, by Theorem 3.1, the least Chebyshev polynomial of degree $n$ becomes
${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})=\left(\frac{1}{W}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(\langle{\mathbf{w}},{\mathbf{x}}\rangle\right).$
We point out that the requirement $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1$, is susceptible to an affine change of variables other than a translation. For instance, the triangle ^$\hat{\triangle}^{2}=\{\,{\mathbf{x}}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\leq 1\,\}$, is a mirror image of $\triangle^{2}$ under $x_{2}\mapsto 1-x_{2}$. However, under the affine change of variables $x_{2}\mapsto 1-x_{2}$, the leading monomial $a_{0}x_{1}^{2}+a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$ of a polynomial in $\Pi_{2}^{*}$ becomes, $a_{0}x_{1}^{2}-a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$, which is no longer an element of $\Pi_{2}^{*}$.
Moreover, the extremal problem (1.3) in Definition 1.1 is not rotationally invariant. To see this consider the following three sets in $\mathbb{R}^{2}$, that are rotations of the same line segment:
$\Omega_{1}$ ::=:= $\{(x,x)\;|\;x\in[-1,1]\}$
$\Omega_{2}$ ::=:= $\{(0,x)\;|\;x\in[-\sqrt{2},\sqrt{2}]\}$
$\Omega_{3}$ ::=:= $\{(x,-x)\;|\;x\in[-1,1]\}.$
Of these three sets, only $\Omega_{1}$ is diagonally-determined. Even though the three sets are rotations of each other, the solution to the extremal problem (1.3) is starkly different in each case:
  • For $\Omega_{1}$, Theorem 3.1 provides the solution $(x_{1},x_{2})\mapsto\frac{1}{2^{n-1}}T_{n}(x_{1})$ as a least Chebyshev polynomial, and thus $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{1}}=\frac{1}{2^{n-1}}$;
  • For $\Omega_{2}$, a least Chebyshev polynomial is clearly given by $(x_{1},x_{2})\mapsto x_{1}^{n}$, so that $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{2}}=0$;
  • For $\Omega_{3}$, a least Chebyshev polynomial is $(x_{1},x_{2})\mapsto\frac{1}{2^{n}}(x_{1}+x_{2})^{n},$ so that $\inf_{P\in\Pi^{*}_{n}}\,\|P\|_{\Omega_{3}}=0$.
The example shows that the diagonal of $\Omega$ plays a key role in determining the least Chebyshev polynomial, and that it is not only an artefact of the analysis.
4. The dual framework and signatures
In this section we consider the dual problem of (1.3). Our goal is to show that, in the special case of diagonally-determined domains, the dual problem has a closed-form solution. Moreover, we will construct a dual solution that is atomic (discrete), and supported on $n+1$ points.
4.1. The dual problem
With $\Omega$ being a domain in ${\mathbb{R}}^{d}$, let $C(\Omega)$ be the space of continuous functions on $\Omega$, and let $C(\Omega)^{*}$ be the dual space of $C(\Omega)$. For compact $\Omega$, one has the following, strong duality result for problem (1.3).
Theorem 4.1.
Let $\Omega\subset{\mathbb{R}}^{d}$ be compact, and consider the dual problem of (1.3), namely
(4.1) $\gamma^{*}\coloneqq\sup_{\begin{subarray}{c}\gamma\in\mathbb{R}\\ L\in C(\Omega)^{*}\end{subarray}}\bigg{\{}\gamma:\;L_{|\Pi_{n-1}^{d}}=0,\,\|L\|_{C(\Omega)^{*}}=1,\mbox{and }L(\mathbf{x}\mapsto\mathbf{x}^{\alpha})=\gamma\;\text{ for all }\alpha\in{\mathbb{N}}^{d}_{n}\bigg{\}},$
where the norm of $L$ is defined as
$\|L\|_{C(\Omega)^{*}}=\sup_{h\in C(\Omega)}\{|Lh|\;:\|h\|_{\Omega}\leq 1\}.$
One has $\gamma^{*}\leq\inf\left\{\|P\|_{\Omega}\;:\;P\in\Pi^{*}_{n}\right\}$ (weak duality), and $\gamma^{*}=\inf\left\{\|P\|_{\Omega}\;:\;P\in\Pi^{*}_{n}\right\}$ if $\Omega$ is compact (strong duality).
One may derive this result directly through conic linear programming duality theory, e.g. [14, Proposition 2.9], and we will omit the proof here.
It is insightful, though, to make a link with classical Chebyshev approximation. To this end, we denote an optimal solution of problem (1.3) by
(4.2) ${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}+Q^{*}({\mathbf{x}})\quad\hbox{such that}\quad\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{{\alpha}}=1,\qquad Q^{*}\in\Pi^{d}_{n-1}.$
Note that $P^{*}\in\Pi_{n}^{*}$. Clearly,
$\|P^{*}\|_{\Omega}=\min_{Q\in\Pi_{n-1}^{d}}\left\|\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}+Q({\mathbf{x}})\right\|_{\Omega},$
while the latter problem is the classical Chebyshev problem of approximating the homogeneous polynomial ${\mathbf{x}}\mapsto\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}$ from $\Pi_{n-1}^{d}$.
Recall the strong duality result for the classical Chebyshev approximation problem as given in the paper by Rivlin and Shapiro [13, Corollary 2]. [^1] In [13, Corollary 2], the objective is in fact given as $\sup|L(f)|$, but the absolute value may be omitted w.l.o.g., since $L$ is feasible for the dual problem if and only if $-L$ is feasible.
Theorem 4.2.
For any $f\in C(\Omega)$, with $\Omega$ compact, one has the following
(4.3) $\min_{Q\in\Pi_{n-1}^{d}}\left\|f-Q\right\|_{\Omega}=\max_{L\in C(\Omega)^{*}}\bigg{\{}L(f):\;L_{|\Pi_{n-1}^{d}}=0,\,\|L\|_{C(\Omega)^{*}}=1\bigg{\}}.$
From this, we can immediately deduce the following relation between the two dual problems.
Lemma 4.3.
For $f({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}$, any optimal solution of problem (4.1) is also optimal for the dual problem in (4.3).
Proof.
Let $L^{*}$ denote an optimal solution of problem (4.1). Then $L^{*}$ is feasible for the dual (maximization) problem in (4.3). Moreover,
$L^{*}(f)=L^{*}\left(\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}\right)=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}L^{*}({\mathbf{x}}^{{\alpha}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}\gamma^{*}=\gamma^{*}=\|P^{*}\|_{\Omega}.$
This yields the result. ∎
In the next subsection, we review the fact that the optimal dual solutions may be assumed to be atomic (discrete) without loss of generality.
4.2. Atomic dual solutions and signatures
We first recall a classical interpolation formula for linear functionals, as given in [13, Corollary 3].
Theorem 4.4.
Let $L$ be any linear functional on a finite dimensional subspace $V$ of $C(\Omega)$, with $\Omega$ compact. Then there exist points $\omega_{1},\ldots,\omega_{r}\in\Omega$ with $r\leq\mathrm{dim}(V)$, and non-zero scalars $\tau_{1},\ldots,\tau_{r}$, such that, defining the point evaluation functionals
$L_{\omega_{i}}(f)=f(\omega_{i})\quad i=1,\ldots,r,$
one has
(4.4) $L=\sum_{i=1}^{r}\tau_{i}L_{\omega_{i}},\quad\|L\|=\sum_{i=1}^{r}|\tau_{i}|.$
As a consequence, there exists a discrete (atomic) solution to problem (4.1). We give a proof below only for the sake of completeness and later reference — the type of argument we use is classical.
Corollary 4.5.
There exist points $\omega_{1},\ldots,\omega_{r}\in\Omega$ with $r\leq\mathrm{dim}(\Pi_{n}^{d})$, and non-zero scalars $\tau_{1},\ldots,\tau_{r}$ with $\sum_{i=1}^{r}|\tau_{i}|=1$, such that an optimal solution of problem (4.1) is given by (4.4). Moreover, the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3), i.e.,
$|P^{*}(\omega_{i})|=\|P^{*}\|_{\Omega}\quad\text{ for all }i=1,\ldots,r.$
Proof.
The first statement follows immediately from Theorem 4.4, applied to an optimal solution of problem (4.1), and using $V=\Pi_{n}^{d}$. It remains to show that the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3). Now let $L^{*}$ denote an optimal solution of problem (4.1), so that
(4.5) $L^{*}=\sum_{i=1}^{r}\tau_{i}L_{\omega_{i}},\quad\sum_{i=1}^{r}|\tau_{i}|=1.$
By Lemma 4.3, $L^{*}$ is also an optimal solution for the dual problem in (4.3). Thus, for any optimal solution $P^{*}$ of problem (1.3), one has
$\|P^{*}\|_{\Omega}=L^{*}(P^{*})=\sum_{i=1}^{r}\tau_{i}P^{*}(\omega_{i}).$
Since $\sum_{i=1}^{r}|\tau_{i}|=1$, $\|P^{*}\|_{\Omega}$ is a weighted average of the values $P^{*}(\omega_{i})$ with $\tau_{i}>0$ and $-P^{*}(\omega_{i})$ with $\tau_{i}<0$. As a consequence, we have
$\|P^{*}\|_{\Omega}=\left\{\begin{array}[]{cc}\phantom{-}P^{*}(\omega_{i})&\mbox{ if $\tau_{i}>0$,}\\ -P^{*}(\omega_{i})&\mbox{if $\tau_{i}<0$,}\end{array}\right.$
completing the proof. ∎
The extremal points in Corollary 4.5 are usually called the support of an extremal signature, defined along the line of [12, Section 2.2] as follows.
Definition 4.6.
A signature with finite support $S\subset\Omega$ is simply a (partition) function from $S$ to {±$\{\pm 1\}$. A signature $\sigma$ with support $S$ is said to be extremal for a subspace $V\subset C(\Omega)$ if there exist weights $\lambda_{\omega}>0$, $\omega\in S$, such that $\sum_{\omega\in S}\lambda_{\omega}\sigma(\omega)v(\omega)=0$ for all $v\in V$. A signature $\sigma$ with support $S$ is said to be associated with a function $g\in C(\Omega)$ if $S$ is included in the set $\{\omega\in\Omega:|g(\omega)|=\|g\|_{\Omega}\}$ of extremal points of $g$ and if $\sigma(\omega)=\operatorname{sgn}(g(\omega))$ for all $\omega\in S$.
Thus, the result of Corollary 4.5 may be restated as the existence of a signature with support $S=\{\omega_{1},\ldots,\omega_{r}\}$ where $r\leq\mbox{dim}(\Pi_{n}^{d})$. This signature is extremal for $\Pi_{n-1}^{d}$, and associated with every optimal solution of problem (1.3). Therefore, we will simply refer to an optimal signature for problem (1.3). Formally we have the following result.
Proposition 4.7.
Any atomic solution of the dual problem (4.1), say $L^{*}$ of the form (4.5), gives rise to an optimal signature for problem (1.3), by setting:
$S=\{\omega_{1},\ldots,\omega_{r}\},\quad\sigma(\omega_{i})=\operatorname{sgn}(\tau_{i}),\quad\lambda_{i}=|\tau_{i}|,\qquad i=1,\ldots,r.$
This is essentially a reformulation of the well-known characterization of best approximation by polynomials in terms of the extremal signature ([12, Theorem 2.6]) for the dual problem.
4.3. Signatures for diagonally-determined domains
For the special case where $\Omega$ is a diagonally-determined set, we may now infer information about an optimal signature from Theorem 3.1 and Proposition 4.7. The idea of the proof is to construct an atomic optimal solution for the dual problem (4.1).
Theorem 4.8.
Let $\Omega$ be a diagonally-determined set with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$. Then, an optimal signature for problem (1.3) is defined by a set $S$, of $n+1$ points
(4.6) $\omega_{j}=a\mathbf{1}+\frac{b-a}{2}\left(1+\cos\left(\frac{(j-1)\pi}{n}\right)\right){\mathbf{1}}\quad\text{ for }j=1,\ldots,n+1,$
where ${\mathbf{1}}$ again denotes the all-ones vector in $\mathbb{R}^{d}$, together with the partition function $\sigma:S\rightarrow\{-1,1\}$, given by
$\sigma(\omega_{j})=\left\{\begin{array}[]{cc}\phantom{-}1&\mbox{if $j$ is odd}\\ -1&\mbox{if $j$ is even,}\end{array}\right.$
as well as $\lambda_{1}=\lambda_{n+1}=\frac{1}{2}$, and $\lambda_{i}=1$ if $i=2,\ldots,n$.
Proof.
By Theorem 3.1, an optimal solution to problem (1.3) is given by the polynomial $P^{*}$ in (3.2), namely
$P^{*}({\mathbf{x}})=\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(-1+2\,\frac{\langle{\mathbf{v}},{\mathbf{x}}\rangle-a}{b-a}\right).$
Using the well-known fact that the extremal points of $T_{n}$ are given by the Gauss-Lobatto-Chebyshev points $\xi_{j}\coloneqq\cos\left((j-1)\pi/n\right)$ for $j=1,\ldots,n+1$, we have that the points listed in (4.6) are extremal points of $P^{*}$. Indeed, using $\langle{\mathbf{v}},\mathbf{1}\rangle=1$, one has
$\langle{\mathbf{v}},\omega_{j}\rangle=a+\frac{b-a}{2}\left(1+\cos\left(\frac{(j-1)\pi}{n}\right)\right)\quad\text{ for }j=1,\ldots,n+1,$
which is the same as
$-1+2\,\frac{\langle{\mathbf{v}},\omega_{j}\rangle-a}{b-a}=\xi_{j}\text{ for }j=1,\ldots,n+1.$
Next, we will construct an atomic solution of the dual problem (4.1), which will lead to the required optimal signature, by Proposition 4.7.
To this end, define the linear operator
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{*}=\frac{1}{n}\sum_{i=1}^{n+1}(-1)^{i+1}\lambda_{i}L_{\omega_{i}}.}$
We claim that $L^{*}$ is an optimal solution of the dual problem (4.1). We first verify feasibility. We will show that $L^{*}$ vanishes on $\Pi^{d}_{n-1}$, as required, by using the Gauss-Lobatto-Chebyshev quadrature formula, which implies, for $k<n$:
[…]
due to the orthogonality of Chebyshev polynomials. Using $T_{n}(\xi_{j})=(-1)^{1+j}$ yields
(4.7) $L^{*}(T_{k})=\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k}(\xi_{j})=0\quad\mbox{ if $k<n$.}$
Thus $L^{*}$ vanishes on univariate (and separable) polynomials in $\Pi^{d}_{n-1}$. For the general case, it suffices to consider $d=2$ and the monomial $p(x_{1},x_{2})=T_{k_{1}}(x_{1})T_{k_{2}}(x_{2})$, where $k_{1}+k_{2}<n$, and show that $L^{*}(p)=0$. To this end, note that
$L^{*}(p)$ $=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}}(\xi_{j})T_{k_{2}}(\xi_{j})$
$=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}\frac{1}{2}[T_{|k_{1}-k_{2}|}(\xi_{j})+T_{k_{1}+k_{2}}(\xi_{j})]$
$=$ $\frac{1}{2n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{|k_{1}-k_{2}|}(\xi_{j})+\frac{1}{2n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}T_{k_{1}+k_{2}}(\xi_{j})=0,$
using (4.7). The result for general $d$ now follows by induction.
Secondly, we verify that $L^{*}(\mathbf{x}^{\alpha})$ is independent of $\alpha$, when $|\alpha|=n$, say $L^{*}(\mathbf{x}^{\alpha})=\hat{\gamma}$ whenever $\alpha\in{\mathbb{N}}^{d}_{n}$. This follows immediately from the construction of $L^{*}$, since all coordinates of the vectors $\omega_{i}$ are equal for all $i=1,\ldots,n+1$. Finally, to prove optimality, we note that, for $P^{*}$ written in the form (4.2),
$\|P^{*}\|_{\Omega}=L^{*}(P^{*})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}L^{*}(\mathbf{x}^{\alpha})=\hat{\gamma}\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}=\hat{\gamma},$
so, by the weak duality theorem, $L^{*}$ is indeed optimal. The required result now follows from Proposition 4.7. ∎
Remark 4.1.
It is interesting to note that — for diagonally-determined sets — there exists an optimal signature for problem (1.3) with support size $n+1$. One can compare this to the general result for compact $\Omega$ of a signature support size of $\mbox{dim}(\Pi^{d}_{n})={n+d\choose d}$. Remarkably the minimum support size is independent of $d$ in the diagonally-determined case. Moreover, we did not use compactness of $\Omega$ in the proof of Theorem 4.8, whereas compactness is needed for the general bound. Finally, note that the optimal signature in Theorem 4.8 depends on $\Omega$ only through its diagonal $[a,b]$, and is independent of ${\mathbf{v}}$.
5. Detecting the diagonally-determined property
The goal of this section is to provide a numerical scheme to detect whether a given compact basic semi-algebraic set
(5.1) $\Omega\,:=\,\{\,{\mathbf{x}}\in\mathbb{R}^{d}:g_{j}({\mathbf{x}})\geq 0\,,\quad j=1,\ldots,m\,\}$
for some given polynomials $g_{j}$, is diagonally-determined.
The quadratic module of $g_{1},\ldots,g_{m}$, denoted by $\mathcal{Q}(g_{1},\ldots,g_{m})$, is defined as all polynomials of the form
(5.2) $p\,=\,\sigma_{0}+\sigma_{1}\,g_{1}+\cdots+\sigma_{m}\,g_{m}\,$
for some sum-of-squares polynomials $\sigma_{0},\sigma_{1},\ldots,\sigma_{m}$. Clearly, if $p\in\mathcal{Q}(g_{1},\ldots,g_{m})$, then $p({\mathbf{x}})\geq 0$ for all ${\mathbf{x}}\in\Omega$.
Assumption 5.1.
We assume that $\mathcal{Q}(g_{1},\ldots,g_{m})$ satisfies the Archimedean condition, meaning that there exists a $p\in\mathcal{Q}(g_{1},\ldots,g_{m})$ such that
$\left\{{\mathbf{x}}\in\mathbb{R}^{d}\;:\;p({\mathbf{x}})\geq 0\right\}\mbox{ is compact.}$.
Under this assumption, if a polynomial is (strictly) positive on $\Omega$, then it belongs to $\mathcal{Q}(g_{1},\ldots,g_{m})$. We state this result — known as Putinar’s Positivstellensatz — as a theorem for later use.
Theorem 5.2 (Putinar [10]).
Assume $\Omega$ is of the form (5.1) and that Assumption 5.1 holds. If $p\in\Pi^{d}_{n}$ satisfies $p({\mathbf{x}})>0$ for all ${\mathbf{x}}\in\Omega$, then $p\in\mathcal{Q}(g_{1},\ldots,g_{m})$.
We also introduce the concept of the truncated quadratic module of degree $k\in\mathbb{N}_{0}$, denoted by $\mathcal{Q}(g_{1},\ldots,g_{m})_{k}$, as the subset of $\mathcal{Q}(g_{1},\ldots,g_{m})$ where each term on the right-hand-side of (5.2) has degree at most $k$, i.e. deg$(\sigma_{0})\leq k$, and deg$(\sigma_{j}g_{j})\leq k$ for all $j=1,\ldots,m$. The membership problem for $\mathcal{Q}(g_{1},\ldots,g_{m})_{k}$ may be reformulated as a semidefinite programming feasibility problem; see Lasserre [6], and the references therein.
Finding the diagonal $[a,b]$
We first note that one may find the diagonal of $\Omega$ by finding the real roots of the univariate polynomials
$t\mapsto p_{j}(t):=g_{j}(t\mathbf{1}),\quad j=1,\ldots,m.$
Indeed, for each $p_{j}$ we may find the set of all (possibly unbounded) intervals where it is nonnegative, and subsequently take the intersection of all these sets for $j=1,\ldots,m$. If the result is a single interval $[a,b]$, then $\mathrm{diag}(\Omega)=[a,b]$, and we may proceed with the next step, since the first property holds in Definition 2.1.
Once $[a,b]$ is known
We assume that we now know $[a,b]=\mathrm{diag}(\Omega)$, and that $a<0<b$. The latter assumption is without loss of generality, due to Lemma 2.7. Then, to detect whether the second property holds in Definition 2.1, consider the optimization problem:
(5.3) $\rho=\displaystyle\min_{{\mathbf{v}}^{\prime}}\,\{\,|1-\langle{\mathbf{v}}^{\prime},\mathbf{1}\rangle|:\>a\leq\langle{\mathbf{v}}^{\prime},{\mathbf{x}}\rangle\,\leq\,b\quad\,\forall{\mathbf{x}}\in\Omega\,\}\,.$
Then $\Omega$ is diagonally-determined if and only if $\rho=0$. Note that one may in fact omit the absolute value in the objective of (5.3), since ${\mathbf{x}}=b\mathbf{1}\in\Omega$, and $\langle{\mathbf{v}}^{\prime},{\mathbf{x}}\rangle\leq b$ therefore implies $\langle{\mathbf{v}}^{\prime},\mathbf{1}\rangle\leq 1$ for all ${\mathbf{v}}^{\prime}$ that are feasible for (5.3).
We now relax this problem to a sequence of semidefinite programs, using the approach introduced by Lasserre [6]. The key idea is to replace the nonnegativity conditions in (5.3) by the sufficient condition of membership of a truncated quadratic module. To this end, we introduce the two affine functions, for given ${\mathbf{v}}^{\prime}\in\mathbb{R}^{d}$:
${\mathbf{x}}\mapsto b-\langle{\mathbf{v}}^{\prime},{\mathbf{x}}\rangle=\ell_{{\mathbf{v}}^{\prime},b}({\mathbf{x}}),\;{\mathbf{x}}\mapsto\langle{\mathbf{v}}^{\prime},{\mathbf{x}}\rangle-a=\ell_{{\mathbf{v}}^{\prime},a}({\mathbf{x}}),$
to obtain the reformulation of problem (5.3) as
[…]
For every $k\in\mathbb{N}_{0}$, we now define:
(5.4) $\rho_{k}=\min_{{\mathbf{v}}^{\prime}}\left\{1-\langle{\mathbf{v}}^{\prime},\mathbf{1}\rangle\;:\;\ell_{{\mathbf{v}}^{\prime},a},\ell_{{\mathbf{v}}^{\prime},b}\in\mathcal{Q}(g_{1},\ldots,g_{m})_{k}\right\}.$
Importantly, and as already mentioned, we emphasize that the optimization problem (5.4) is a (convex) semidefinite program that can be solved efficiently with arbitrary precision (fixed in advance), under mild assumptions, using, e.g., interior point methods; see Lasserre [6].
Proposition 5.3.
Assume $\Omega$ is of the form (5.1), that Assumption 5.1 holds, and that $\mathrm{diag}(\Omega)=[a,b]$ where $a<0<b$. Then one has $\rho_{k}\in[0,1]$ for all $k$. Moreover, if $\Omega$ is diagonally-determined, then the sequence $\{\rho_{k}\}$ converges monotonically to zero. If $\rho_{k}=0$ for some finite value of $k$, then the optimal ${\mathbf{v}}^{\prime}\in\mathbb{R}^{d}$ of (5.4) certifies that $\Omega$ is diagonally-determined with vector ${\mathbf{v}}={\mathbf{v}}^{\prime}$ and diagonal $[a,b]$.
Proof.
For $k=0$, a feasible solution of (5.4) is given by ${\mathbf{v}}^{\prime}=\mathbf{0}$, since $a<0<b$, so that $\rho_{0}\leq 1$. Consequently, $\rho_{k}\in[0,1]$ for all $k$, since $\mathcal{Q}(g_{1},\ldots,g_{m})_{k}\subseteq\mathcal{Q}(g_{1},\ldots,g_{m})_{k+1}$ for all $k\in\mathbb{N}_{0}$. Assume now that $\Omega$ is diagonally-determined with vector ${\mathbf{v}}$ and diagonal $[a,b]$. Then, for any $\varepsilon>0$ and ${\mathbf{x}}\in\Omega$, one has
$\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle\in[(1-\varepsilon)a,(1-\varepsilon)b]\subset(a,b),$
where we again use that $a<0<b$. Thus, ${\mathbf{x}}\mapsto\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle-a\in\mathcal{Q}(g_{1},\ldots,g_{m})$ and ${\mathbf{x}}\mapsto b-\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle\in\mathcal{Q}(g_{1},\ldots,g_{m})$, by Theorem 5.2. This implies that, for sufficiently large $k^{*}\in\mathbb{N}_{0}$, ${\mathbf{v}}^{\prime}=(1-\varepsilon){\mathbf{v}}$ is feasible for (5.4) with $k=k^{*}$. Thus $\rho_{k^{*}}\leq 1-\langle(1-\varepsilon){\mathbf{v}},\mathbf{1}\rangle=\varepsilon$, and therefore the sequence $\{\rho_{k}\}$ converges monotonically to zero. ∎
Example 5.4.
It is insightful to see how the above procedure works in the case of the unit ball in the $\infty$-norm, i.e.
$\Omega=\{{\mathbf{x}}\in\mathbb{R}^{d}\;:\;\|{\mathbf{x}}\|_{\infty}\leq 1\}.$
By Proposition 2.3, and its proof, we know that $\Omega$ is diagonally-determined, with diagonal $[a,b]=[-1,1]$ and vector ${\mathbf{v}}=\frac{1}{d}\mathbf{1}$. We may describe $\Omega$ as a semi-algebraic set in the form (5.1) by using the polynomials $g_{j}({\mathbf{x}})=1-x_{j}^{2}$ for $j=1,\ldots,d$, i.e. $m=d$. The next step is to find the diagonal of $\Omega$ by finding the roots of the univariate polynomials $p_{j}(t):=g_{j}(t\mathbf{1})=1-t^{2}$ for $j=1,\ldots,d$. These polynomials are the same for all $j$ and nonnegative on the interval $[-1,1]$, so that we obtain $[a,b]=[-1,1]$. The next step is to solve (5.4) for increasing values of $k\in\mathbb{N}$. Using the identity
$1\pm x=\frac{1}{2}\left((1\pm x)^{2}+1-x^{2}\right),\quad(x\in\mathbb{R}),$
one obtains that
$b-\langle{\mathbf{v}},{\mathbf{x}}\rangle=1-\frac{1}{d}\sum_{j=1}^{d}x_{j}=\frac{1}{d}\sum_{j=1}^{d}\left(1-x_{j}\right)=\frac{1}{2d}\sum_{j=1}^{d}\left((1-x_{j})^{2}+g_{j}({\mathbf{x}})\right).$
Similarly one has
$\langle{\mathbf{v}},{\mathbf{x}}\rangle-a=\frac{1}{2d}\sum_{j=1}^{d}\left((1+x_{j})^{2}+g_{j}({\mathbf{x}})\right).$
Thus, for $k=2$, we find that ${\mathbf{v}}^{\prime}={\mathbf{v}}$ is feasible in (5.4), so that $\rho_{2}=0$.
In the example we have finite convergence of the sequence $\{\rho_{k}\}$, but it is an open question when this occurs. There are known sufficient conditions for the finite convergence of Lasserre-type semidefinite programming hierarchies, e.g. [8], but these are not satisfied for our problem in general.
6. Conclusion and discussion
[…]
${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})=\left(\frac{1}{\|\mathbf{1}\|}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(\|\mathbf{1}\|\langle{\mathbf{v}},{\mathbf{x}}\rangle\right).$
Since $\frac{1}{2^{n-1}}T_{n}$ is a monic univariate polynomial of degree $n$, the leading coefficients of the least Chebyshev polynomial $P^{*}$ are therefore obtained from the multinomial expansion of ${\mathbf{x}}\mapsto\langle{\mathbf{v}},{\mathbf{x}}\rangle^{n}$. Since ${\mathbf{v}}$ is entrywise-nonnegative by assumption, it follows that these leading coefficients are also nonnegative. In this sense our generalization of monic polynomials to the multivariate case is a natural one, but is remains an interesting question which diagonally-determined sets allow a entrywise-nonnegative vector ${\mathbf{v}}$.
A second point to mention, is that we have not explored the geometric and invariance properties of diagonally-determined domains to any great extent in this paper, and this deserves to be investigated further. The same may be said for the decision problem of whether a given domain is diagonally-determined. If $\Omega$ is a compact semi-algebraic set, one may formulate an algorithmic procedure to decide this question, as we did in Section 5. An interesting special case is the (complexity) question: may one decide in polynomial time is a given polyhedron is diagonally-determined?
Finally, we note that our main results in Theorems 3.1 and 4.8 are relatively straightforward to prove, once the correct formulations have been discovered. We ‘discovered’ these results by doing numerical experiments on least Chebyshev polynomials for the Euclidean unit ball, using the tools described in our earlier paper [4]. Thus we observed the structure of the atomic dual solution in Theorem 4.8, and we could infer the main results from there. In the immortal words of Bernhard Riemann: [^2] As quoted in I. Lakatos, Proofs and Refutations (Cambridge University Press, 1976).
If only I had the theorems! Then I should find the proofs easily enough.
Acknowledgments
This work is the result of a collaboration made possible by the SQuaRE program at the American Institute of Mathematics (AIM). We are truly grateful to AIM for the supportive and mathematically rich environment they provided. In addition, we owe thanks to several funding agencies, as M. D. is supported by the Australian Research Council Discovery Early Career Award DE240100674, S. F. is partially supported by grants from the National Science Foundation (DMS-2053172) and from the Office of Naval Research (N00014-20-1-2787), E. de K. is supported by grants from the Dutch Research Council (NWO) (OCENW.M.23.050 and OCENW.GROOT.2019.015), J. B. L. is supported by the AI Interdisciplinary Institute through the French program “Investing for the Future PI3A” (ANR-19-PI3A-0004) and by the National Research Foundation, Singapore, through the DesCartes and Campus for Research Excellence and Technological Enterprise (CREATE) programs, and Y. X. is partially supported by the Simons Foundation (grant #849676).